OpenStudy (anonymous):

What is the solution set of 4x^2 - 36 = 0

OpenStudy (vishweshshrimali5):

\[4x^2 - 36 = 0 => x^2 - 9 = 0 => (x+3)(x-3) = 0 => x = -3 or 3\]

OpenStudy (jackellyn):

4(x^2 - 9)=0 4(x+3)(x-3)=0 x=-3 or x=3

OpenStudy (anonymous):

+- 3